4 ms·
The article gets the 50/50 case subtly wrong. The really naive analysis is that you have two doors so it's 50/50. A less naive, more interesting analysis, is th
by pdpi 4y ago
The article gets the 50/50 case subtly wrong. The really naive analysis is that you have two doors so it's 50/50. A less naive, more interesting analysis, is that there are twelve possible outcomes, of which six are favourable (if the correct option is A, pick A -> presenter picks B is different from pick A -> presenter picks C).
The article hides this away by lumping those two under "Host opens B or C" without further justification, but it's important to notice that this only works because those twelve outcomes have different probabilities.
Edit: In table form,
Car Guest Monty Swap? Probability
A A B No 1/3 * 1/3 * 1/2 = 1/18
A A C No 1/3 * 1/3 * 1/2 = 1/18
A B C Yes 1/3 * 1/3 = 1/9
A C B Yes 1/3 * 1/3 = 1/9
B A C Yes 1/3 * 1/3 = 1/9
B B A No 1/3 * 1/3 * 1/2 = 1/18
B B C No 1/3 * 1/3 * 1/2 = 1/18
B C A Yes 1/3 * 1/3 = 1/9
C A B Yes 1/3 * 1/3 = 1/9
C B A Yes 1/3 * 1/3 = 1/9
C C A No 1/3 * 1/3 * 1/2 = 1/18
C C B No 1/3 * 1/3 * 1/2 = 1/18
Total Yes = 6 * 1/9 = 6/9 = 2/3
Total No = 6 * 1/18 = 6/18 = 1/3